Find the slope of the line through each pair of points
(5,-4),(-20,-3)

Answer by ewatrrr(10682)

 
Hi,
Find the slope of the line through each pair of points:


(5,-4),
(-20,-3) m = -1/25
y = (-1/25)x + b ||Using Pt(5,-4) to solve for b
-4 = (-1/25)5 + b
-19/5 = b
y = (-1/25)x -19/5 or in standard Form: x + 25y = -95